OpenStudy (edwin99):

At the beginning of every year, Molly deposits $200 in a savings account that offers an interest rate of 20%, compounded annually. The total amount that Molly will have in her account at the end of 3 years is $

OpenStudy (rolldextro):

SImple Interest Formula \[FV = PV (1+i n) |\] =440

OpenStudy (rolldextro):

Fv=200)(1+6*.20)